# Tree Embedding in High Dimensions:

### v1

@@ 1.2.2 Proof Overview: Implementations in $\mathbb{R}^{d}$ and Applications
We employ consistent hashing [CJK+22, FJL+25] techniques to utilize the locality property of our tree embedding algorithm in $\mathbb{R}^{d}$ (see Definition˜4.1). Roughly speaking, such hashing $\varphi:\mathbb{R}^{d}\to U$ yields data-oblivious $\tau$-bounded partitions (for any $\tau$) such that for any $B(p,\tau/\Gamma)$ it holds that $|\varphi(B(p,\tau/\Gamma))|\leq\Lambda:=n^{\operatorname{poly}(1/\Gamma)}$ for any $p\in\mathbb{R}^{d}$. This exactly means $|\tilde{B}(p,r)|\leq\Lambda=n^{\operatorname{poly}(1/\Gamma)}$ as in our tree embedding algorithm. Crucially, for the purposes of our dynamic implementation, the hash function of [CJK+22, FJL+25] allows for the computation of the value $\varphi(p)$ and set $\varphi(B(p,\tau/\Gamma))$ in $\operatorname{poly}(d)$ and $|\varphi(B(p,\tau/\Gamma))|\cdot\operatorname{poly}(d)$ time respectively.
@@
The MPC implementation follows immediately by utilizing a low-space construction of consistent hashing [CJK+22] (see Lemma˜4.5). In the dynamic setting, we need to plug in a time-efficient consistent hashing bound, albeit the parameter is slightly worse (see Lemma˜4.4). However, the dynamic implementation requires even more work. The first observation is that deletions can be “ignored”, since deleting a data point may be interpreted as deactivating a leaf node in an existing embedding tree, and doing so does not hurt the distortion. Now suppose we work with insertion-only setting. Consider an insertion of a point $p$ and some scale $w_{i}$. We first insert $p$ to its bucket under $\varphi_{i}$ and maintain the min $\pi$-value of the bucket. The bucket $\operatorname{buks}(p)$ may intersect many $B(q,r_{i})$’s (over $q\in P$), and we need to scan and update the min $\pi$-value for all such $q$. To do this efficiently, for each $q$ we maintain a priority queue for the intersecting buckets $\varphi_{i}(B(q,r_{i}))$, with respect to the (min) $\pi$-value of the bucket, so the total scan caused by inserting $p$ is proportional to the change of the min element of the priority queues. The total/amortized update time can be bounded using two facts: a) each $|\varphi_{i}(B(q,r_{i}))|\leq\Lambda$ which follows from consistent hashing; and b) the change of min for all priority queues combined is small, which is argued via a low-recourse fact of permutations/random maps, also employed in [MS09]: let $\pi$ be a permutation of $[n]$, then the local minimum $\mathbb{E}[\{i:\pi(i)=\min\{\pi(j):j<i\}\}]=O(\log n)$.
@@
The high-level idea of our applications is to utilize the simple structure of tree embedding, i.e., HST, to find/maintain optimal solutions on the tree. The applications in MPC setting is relatively straightforward, since previous results [CALNF+21, AAH+23] works with a general tree embedding in MPC in a black-box way. Our dynamic applications generally do not work in a black-box way, and we need to introduce new steps to maintain the optimal solution on a tree. We crucially exploit that the dynamic tree embedding maintained by our framework only undergoes $\tilde{O}(1)$ simple changes per updates to the underlying input points. Namely, the embedding itself may only undergo leaf deletions and the insertion of leafs with a path connecting them to existing nodes (for a precise description see theorem˜4.2).
@@ 2 Preliminaries
For integer $n$, let $n]:=\{1,\ldots,n\}$, and let $H_{n}:=\sum_{i\in[n]}1/i$ be the harmonic sum. For a function $\varphi:X\to Y$, the image of $S\subseteq X$ is defined $\varphi(S):=\{\varphi(x):x\in S\}$, and the preimage of $y\subseteq Y$ is defined $\varphi^{-1}(y):=\{x\in X:\varphi(x)=y\}$. Let $(V,\operatorname{dist})$ be an underlying metric space. In Section˜3 we use this general metric space notation, while in later sections we assume it is Euclidean $\mathbb{R}^{d}$, i.e., $V=\mathbb{R}^{d}$ and $\operatorname{dist}=\ell_{2}$. We use $\operatorname{dist}_{\mathcal{T}}(x,y)$ to denote the distance between points $x$ and $y$ in a graph $\mathcal{T}$, which is supposed to be the embedding tree. Let $B(p,r):=\{q:\operatorname{dist}(p,q)\leq r\}$ be the metric ball. For a point set $S$, let $\operatorname{diam}(S):=\max_{x,y\in S}\operatorname{dist}(x,y)$ be the diameter.
@@ 3 Tree Embedding via Metric Decomposition
In this section we consider the general metric space setting and present our framework for tree embedding, listed in Algorithm˜1. Recall that $(V,\operatorname{dist})$ is the notation for the underlying metric space. This algorithm receives as input a dataset $P\subseteq V$, a parameter $\Gamma\geq 1$, and assumes black-box access to a series of $m$ generic metric hash $\varphi_{1},\ldots,\varphi_{m}$, such that for each $i\in[m]$, $\varphi_{i}:V\to U$ for some domain $U$ satisfies that for each image $z\in\varphi(V)$, the diameter of the preimage (which may be viewed as a hash bucket which we define as a notation later) is bounded by $\tau_{i}$, i.e., $\operatorname{diam}(\varphi^{-1}(z))\leq\tau_{i}$ where $\tau_{i}:=w_{i}/2$. The technical core of this algorithm is a modified CKR decomposition [CKR04] operated on general metric hashing, where the major new step is Line 1. As we mentioned in Section˜1.2, this new Line 1 introduces nontrivial technical challenges which requires new ideas in the analysis.
This algorithm serves as an outline for our dynamic and MPC algorithms, and in this section, we only focus on the distortion analysis, i.e., Lemma˜3.1, since the efficiency (in terms of time/round complexity etc.) depends on the actual dynamic/MPC implementations which will be discussed in later sections. We denote by $\mathcal{T}$ the (implicit) embedding tree generated by our algorithm.
@@
Instead of explicitly building an embedding tree $\mathcal{T}$, the output of the algorithm is a sequence of labels $(\ell_{p}^{(1)},\ldots,\ell_{p}^{(m)})$ for each $p\in P$. This is slightly different to classical constructions [FRT04], where the label is built in a top-down recursive manner (whereas our labels are built independently at every level). Our version is similar to that used in previous parallel algorithms for tree embedding [GL14]. To build the tree $\mathcal{T}$ implicitly, for $p\in P$, its labels $(\ell^{(1)}_{p},\ldots,\ell^{(m)}_{p})$ identify each tree nodes from leaf $p$ to the root, as follows: for $i\in[m]$, the $i$-prefix $(\ell^{(1)}_{p},\ldots,\ell^{(i)}_{p})$ constitutes the ID of the level-$i$ ancestor of leaf $p$. The edge length from a level-$i$ node to level-$(i+1)$ node is $w_{i}$, as we discussed in Section˜2.
Notations for Algorithm˜1.
@@ 7 $r_{i}\leftarrow\beta/\Gamma\cdot w_{i}$
The main result of this section establishes that Algorithm˜1 computes a tree embedding achieving near-optimal distortion, as stated below.
@@
Consider an underlying metric $(V,\operatorname{dist})$. Suppose Algorithm˜1 takes a point set $P\subseteq V$ and parameter $\Gamma\geq 1$ as input, and is given access to a series of metric hash $\varphi_{1},\ldots,\varphi_{m}$ each with diameter bound $\tau_{i}:=w_{i}/2$, Let $\mathcal{T}$ be the corresponding embedding tree with respect to the output labels of Algorithm˜1, then $\forall p,q\in P$, it holds that $\operatorname{dist}_{\mathcal{T}}(p,q)\geq\operatorname{dist}(p,q)$ and $\mathbb{E}[\operatorname{dist}_{\mathcal{T}}(p,q)]\leq O(\Gamma\log\Gamma)\cdot\log n\cdot\operatorname{dist}(x,y)$.
@@
The proof can be found in Section˜3.1. ∎
3.1 Proof of Lemma˜3.1: Distortion Analysis of Algorithm˜1
In this section, we prove that the embedding tree corresponding to the output of Algorithm˜1 achieves the desired distortion guarantees. To this end, we begin by defining the distance between two points $p,q\in P$ in $\mathcal{T}$ in terms of their corresponding labels.
@@
    This is achieved by applying the key lemma Lemma˜3.3.
@@
    using a weaker bound Lemma˜3.5.
@@
We note that the proofs of Lemma˜3.3 and its consequence, Lemma˜3.5, are deferred to Section˜3.2. We now proceed to discuss these three cases in more detail.
Analysis of Small-distance case (Item˜3): $i\in[i_{\max}+1,m]$.
@@
Analysis of Large-distance case (Item˜1): $i\in[1,i_{\min}]$ .
@@
The proof can be found in Section˜3.2. ∎
Applying Lemma˜3.3 with $r^{\prime}=r_{i+2}^{\max}$ (which satisfies the condition because of (6)), we obtain
@@
$B(p,r_{i}^{\max})\subseteq\tilde{B}_{j}^{P}(p,r_{i}^{\max})$ follows from the definition of $\tilde{B}_{j}^{P}$. It remains to show $\tilde{B}_{i+t}^{P}(p,r_{i+t}^{\max})\subseteq B(p,r_{i}^{\max})$. Consider an arbitrary $q\in\tilde{B}_{i+t}^{P}(p,r_{i+t}^{\max})$ and we need to show $q\in B(p,r_{i}^{\max})$. Recall that $\tilde{B}_{i+t}^{P}(p,r_{i+t}^{\max})=\operatorname{buks}^{P}_{i+t}(B(p,r_{i+t}^{\max}))$, then there exists some $p^{\prime}\in B(p,r_{i+t}^{\max})$ such that $\varphi_{i+t}(p^{\prime})=\varphi_{i+t}(q)$. By the diameter guarantee of $\varphi$ (see Definition˜4.1), we have $\operatorname{dist}(p,q)\leq\operatorname{dist}(p,p^{\prime})+\operatorname{dist}(p^{\prime},q)\leq r_{i+t}^{\max}+\tau_{i+t}=(1+2\Gamma)r_{i+t}^{\max}\leq 4\Gamma\cdot r_{i+t}^{\max}\leq r_{i}^{\max}$. This finishes the proof of Lemma˜3.4. ∎
Apply Lemma˜3.4 on (7) and let $t:=\lceil\log_{2}\Gamma\rceil+2$ be as in the lemma,
@@
concluding the large-distance case (Item˜1).
Analysis of Similar-distance case (Item˜2): $i\in[i_{\min}+1,i_{\max}]$.
Following a similar overall strategy with the previous case, we need the following lemma as a universal upper bound of the event $\ell_{p}^{(i)}\neq\ell_{q}^{(i)}$, without any other assumption on $p$ and $q$ (e.g., $B(p,r^{\prime})\subseteq B(q,\frac{1}{2}r_{i}^{\max})$) as required in Lemma˜3.3.
@@
The proof can be found in Section˜3.2. ∎
We conclude the similar-distane case by applying lemma˜3.5
@@
Finally, we are ready to prove Lemma˜3.1
Concluding Lemma˜3.1.
@@
bounding the expected expansion of the tree embedding. This finishes the proof of Lemma˜3.1. $\square$
3.2 Proof of Lemma˜3.3 and Lemma˜3.5
In this section, we focus on proving the two key lemmas, Lemma˜3.3 and Lemma˜3.5, which are used in the proof of Lemma˜3.1 in the analysis of the large-distance (Item˜1) and similar-distance (Item˜2) cases, respectively. The proof of the latter lemma, Lemma˜3.5, follows closely from that of Lemma˜3.3 and is briefly outlined at the end.
@@
Before presenting the proof of Lemma˜3.3, we introduce two key concepts that help us overcome these difficulties, namely the representative set and the key representative.
@@
Our proof strategy is to utilize the structure of $\tilde{B}_{i}^{P}(\cdot,\cdot)$, and argue that a carefully chosen set of representatives of $\tilde{B}_{i}^{P}(p,r)$ has similar geometry properties as a standard ball $B(p,r)$. Specifically, since $\tilde{B}_{i}^{P}$ is a union of buckets, our plan is to pick from each bucket exactly one point define a representative set for $\tilde{B}_{i}^{P}(p,r)$. This representative set aligns better with the geometry of $B(p,r)$, and we formally define the following properties that a proper representative set should satisfy. We show in Lemma˜3.7 that such proper representative set always exists. Then in the remainder of the proof, we work with this representative set instead of directly with $\tilde{B}_{i}^{P}$. While such representative set reveals the major structure of $\tilde{B}_{i}^{P}$, we note that it is only used in the analysis and the algorithm does not rely on it.
@@
Now we turn to showing that such representative sets exist. A natural attempt to define $\operatorname{rep}(\cdot,\cdot)$ such that it satisfies all three of distinct, monotone and ball-preserving properties, is to let $\operatorname{rep}_{i}(r,p):=B(p,r)\cap\operatorname{buks}_{i}(P)$, which is precisely the first half of the property Item˜3. The issue is that it does not satisfy the distinctness Item˜1. Our definition presented in the Lemma˜3.7 below fixes this by only selecting a subset of $B(p,r)\cap\operatorname{buks}_{i}(P)$, roughly the nearest neighbors from $p$ in each bucket of $B(p,r)\cap\operatorname{buks}_{i}(P)$. This “nearest” property ensures the distinctness (Item˜1). Using the geometric properties of the nearest neighbor, we can also show that such sets satisfy also properties Item˜2 and Item˜3.
@@
There exists $\{\operatorname{rep}_{i}(p,r):i\in[m],p\in P,r\geq 0\}$ that satisfies Definition˜3.6.

### v2

@@ 1.2.2 Proof Overview: Implementations in $\mathbb{R}^{d}$ and Applications
We employ consistent hashing [CJK+22, FJL+25] techniques to utilize the locality property of our tree embedding algorithm in $\mathbb{R}^{d}$ (see Definition 4.1). Roughly speaking, such hashing $\varphi:\mathbb{R}^{d}\to U$ yields data-oblivious $\tau$-bounded partitions (for any $\tau$) such that for any $B(p,\tau/\Gamma)$ it holds that $|\varphi(B(p,\tau/\Gamma))|\leq\Lambda:=n^{\operatorname{poly}(1/\Gamma)}$ for any $p\in\mathbb{R}^{d}$. This exactly means $|\tilde{B}(p,r)|\leq\Lambda=n^{\operatorname{poly}(1/\Gamma)}$ as in our tree embedding algorithm. Crucially, for the purposes of our dynamic implementation, the hash function of [CJK+22, FJL+25] allows for the computation of the value $\varphi(p)$ and set $\varphi(B(p,\tau/\Gamma))$ in $\operatorname{poly}(d)$ and $|\varphi(B(p,\tau/\Gamma))|\cdot\operatorname{poly}(d)$ time respectively.
@@
The MPC implementation follows immediately by utilizing a low-space construction of consistent hashing [CJK+22] (see Lemma 4.5). In the dynamic setting, we need to plug in a time-efficient consistent hashing bound, albeit the parameter is slightly worse (see Lemma 4.4). However, the dynamic implementation requires even more work. The first observation is that deletions can be “ignored”, since deleting a data point may be interpreted as deactivating a leaf node in an existing embedding tree, and doing so does not hurt the distortion. Now suppose we work with insertion-only setting. Consider an insertion of a point $p$ and some scale $w_{i}$. We first insert $p$ to its bucket under $\varphi_{i}$ and maintain the min $\pi$-value of the bucket. The bucket $\operatorname{buks}(p)$ may intersect many $B(q,r_{i})$’s (over $q\in P$), and we need to scan and update the min $\pi$-value for all such $q$. To do this efficiently, for each $q$ we maintain a priority queue for the intersecting buckets $\varphi_{i}(B(q,r_{i}))$, with respect to the (min) $\pi$-value of the bucket, so the total scan caused by inserting $p$ is proportional to the change of the min element of the priority queues. The total/amortized update time can be bounded using two facts: a) each $|\varphi_{i}(B(q,r_{i}))|\leq\Lambda$ which follows from consistent hashing; and b) the change of min for all priority queues combined is small, which is argued via a low-recourse fact of permutations/random maps, also employed in [MS09]: let $\pi$ be a permutation of $[n]$, then the local minimum $\mathbb{E}[\{i:\pi(i)=\min\{\pi(j):j<i\}\}]=O(\log n)$.
@@
The high-level idea of our applications is to utilize the simple structure of tree embedding, i.e., HST, to find/maintain optimal solutions on the tree. The applications in MPC setting is relatively straightforward, since previous results [CALNF+21, AAH+23] works with a general tree embedding in MPC in a black-box way. Our dynamic applications generally do not work in a black-box way, and we need to introduce new steps to maintain the optimal solution on a tree. We crucially exploit that the dynamic tree embedding maintained by our framework only undergoes $\tilde{O}(1)$ simple changes per updates to the underlying input points. Namely, the embedding itself may only undergo leaf deletions and the insertion of leafs with a path connecting them to existing nodes (for a precise description see Theorem 4.2).
@@ 2 Preliminaries
For integer $n$, let $n]:=\{1,\ldots,n\}$, and let $H_{n}:=\sum_{i\in[n]}1/i$ be the harmonic sum. For a function $\varphi:X\to Y$, the image of $S\subseteq X$ is defined $\varphi(S):=\{\varphi(x):x\in S\}$, and the preimage of $y\subseteq Y$ is defined $\varphi^{-1}(y):=\{x\in X:\varphi(x)=y\}$. Let $(V,\operatorname{dist})$ be an underlying metric space. In Section 3 we use this general metric space notation, while in later sections we assume it is Euclidean $\mathbb{R}^{d}$, i.e., $V=\mathbb{R}^{d}$ and $\operatorname{dist}=\ell_{2}$. We use $\operatorname{dist}_{\mathcal{T}}(x,y)$ to denote the distance between points $x$ and $y$ in a graph $\mathcal{T}$, which is supposed to be the embedding tree. Let $B(p,r):=\{q:\operatorname{dist}(p,q)\leq r\}$ be the metric ball. For a point set $S$, let $\operatorname{diam}(S):=\max_{x,y\in S}\operatorname{dist}(x,y)$ be the diameter.
@@ 3 Tree Embedding via Metric Decomposition
In this section we consider the general metric space setting and present our framework for tree embedding, listed in Algorithm 1. Recall that $(V,\operatorname{dist})$ is the notation for the underlying metric space. This algorithm receives as input a dataset $P\subseteq V$, a parameter $\Gamma\geq 1$, and assumes black-box access to a series of $m$ generic metric hash $\varphi_{1},\ldots,\varphi_{m}$, such that for each $i\in[m]$, $\varphi_{i}:V\to U$ for some domain $U$ satisfies that for each image $z\in\varphi(V)$, the diameter of the preimage (which may be viewed as a hash bucket which we define as a notation later) is bounded by $\tau_{i}$, i.e., $\operatorname{diam}(\varphi^{-1}(z))\leq\tau_{i}$ where $\tau_{i}:=w_{i}/2$. The technical core of this algorithm is a modified CKR decomposition [CKR04] operated on general metric hashing, where the major new step is Line 1. As we mentioned in Section 1.2, this new Line 1 introduces nontrivial technical challenges which requires new ideas in the analysis.
This algorithm serves as an outline for our dynamic and MPC algorithms, and in this section, we only focus on the distortion analysis, i.e., Lemma 3.1, since the efficiency (in terms of time/round complexity etc.) depends on the actual dynamic/MPC implementations which will be discussed in later sections. We denote by $\mathcal{T}$ the (implicit) embedding tree generated by our algorithm.
@@
Instead of explicitly building an embedding tree $\mathcal{T}$, the output of the algorithm is a sequence of labels $(\ell_{p}^{(1)},\ldots,\ell_{p}^{(m)})$ for each $p\in P$. This is slightly different to classical constructions [FRT04], where the label is built in a top-down recursive manner (whereas our labels are built independently at every level). Our version is similar to that used in previous parallel algorithms for tree embedding [GL14]. To build the tree $\mathcal{T}$ implicitly, for $p\in P$, its labels $(\ell^{(1)}_{p},\ldots,\ell^{(m)}_{p})$ identify each tree nodes from leaf $p$ to the root, as follows: for $i\in[m]$, the $i$-prefix $(\ell^{(1)}_{p},\ldots,\ell^{(i)}_{p})$ constitutes the ID of the level-$i$ ancestor of leaf $p$. The edge length from a level-$i$ node to level-$(i+1)$ node is $w_{i}$, as we discussed in Section 2.
Notations for Algorithm 1.
@@ 7 $r_{i}\leftarrow\beta/\Gamma\cdot w_{i}$
The main result of this section establishes that Algorithm 1 computes a tree embedding achieving near-optimal distortion, as stated below.
@@
Consider an underlying metric $(V,\operatorname{dist})$. Suppose Algorithm 1 takes a point set $P\subseteq V$ and parameter $\Gamma\geq 1$ as input, and is given access to a series of metric hash $\varphi_{1},\ldots,\varphi_{m}$ each with diameter bound $\tau_{i}:=w_{i}/2$, Let $\mathcal{T}$ be the corresponding embedding tree with respect to the output labels of Algorithm 1, then $\forall p,q\in P$, it holds that $\operatorname{dist}_{\mathcal{T}}(p,q)\geq\operatorname{dist}(p,q)$ and $\mathbb{E}[\operatorname{dist}_{\mathcal{T}}(p,q)]\leq O(\Gamma\log\Gamma)\cdot\log n\cdot\operatorname{dist}(x,y)$.
@@
The proof can be found in Section 3.1. ∎
3.1 Proof of Lemma 3.1: Distortion Analysis of Algorithm 1
In this section, we prove that the embedding tree corresponding to the output of Algorithm 1 achieves the desired distortion guarantees. To this end, we begin by defining the distance between two points $p,q\in P$ in $\mathcal{T}$ in terms of their corresponding labels.
@@
    This is achieved by applying the key lemma Lemma 3.3.
@@
    using a weaker bound Lemma 3.5.
@@
We note that the proofs of Lemma 3.3 and its consequence, Lemma 3.5, are deferred to Section 3.2. We now proceed to discuss these three cases in more detail.
Analysis of Small-distance case (Item 3): $i\in[i_{\max}+1,m]$.
@@
Analysis of Large-distance case (Item 1): $i\in[1,i_{\min}]$ .
@@
The proof can be found in Section 3.2. ∎
Applying Lemma 3.3 with $r^{\prime}=r_{i+2}^{\max}$ (which satisfies the condition because of (6)), we obtain
@@
$B(p,r_{i}^{\max})\subseteq\tilde{B}_{j}^{P}(p,r_{i}^{\max})$ follows from the definition of $\tilde{B}_{j}^{P}$. It remains to show $\tilde{B}_{i+t}^{P}(p,r_{i+t}^{\max})\subseteq B(p,r_{i}^{\max})$. Consider an arbitrary $q\in\tilde{B}_{i+t}^{P}(p,r_{i+t}^{\max})$ and we need to show $q\in B(p,r_{i}^{\max})$. Recall that $\tilde{B}_{i+t}^{P}(p,r_{i+t}^{\max})=\operatorname{buks}^{P}_{i+t}(B(p,r_{i+t}^{\max}))$, then there exists some $p^{\prime}\in B(p,r_{i+t}^{\max})$ such that $\varphi_{i+t}(p^{\prime})=\varphi_{i+t}(q)$. By the diameter guarantee of $\varphi$ (see Definition 4.1), we have $\operatorname{dist}(p,q)\leq\operatorname{dist}(p,p^{\prime})+\operatorname{dist}(p^{\prime},q)\leq r_{i+t}^{\max}+\tau_{i+t}=(1+2\Gamma)r_{i+t}^{\max}\leq 4\Gamma\cdot r_{i+t}^{\max}\leq r_{i}^{\max}$. This finishes the proof of Lemma 3.4. ∎
Apply Lemma 3.4 on (7) and let $t:=\lceil\log_{2}\Gamma\rceil+2$ be as in the lemma,
@@
concluding the large-distance case (Item 1).
Analysis of Similar-distance case (Item 2): $i\in[i_{\min}+1,i_{\max}]$.
Following a similar overall strategy with the previous case, we need the following lemma as a universal upper bound of the event $\ell_{p}^{(i)}\neq\ell_{q}^{(i)}$, without any other assumption on $p$ and $q$ (e.g., $B(p,r^{\prime})\subseteq B(q,\frac{1}{2}r_{i}^{\max})$) as required in Lemma 3.3.
@@
The proof can be found in Section 3.2. ∎
We conclude the similar-distane case by applying Lemma 3.5
@@
Finally, we are ready to prove Lemma 3.1
Concluding Lemma 3.1.
@@
bounding the expected expansion of the tree embedding. This finishes the proof of Lemma 3.1. ∎
3.2 Proof of Lemma 3.3 and Lemma 3.5
In this section, we focus on proving the two key lemmas, Lemma 3.3 and Lemma 3.5, which are used in the proof of Lemma 3.1 in the analysis of the large-distance (Item 1) and similar-distance (Item 2) cases, respectively. The proof of the latter lemma, Lemma 3.5, follows closely from that of Lemma 3.3 and is briefly outlined at the end.
@@
Before presenting the proof of Lemma 3.3, we introduce two key concepts that help us overcome these difficulties, namely the representative set and the key representative.
@@
Our proof strategy is to utilize the structure of $\tilde{B}_{i}^{P}(\cdot,\cdot)$, and argue that a carefully chosen set of representatives of $\tilde{B}_{i}^{P}(p,r)$ has similar geometry properties as a standard ball $B(p,r)$. Specifically, since $\tilde{B}_{i}^{P}$ is a union of buckets, our plan is to pick from each bucket exactly one point define a representative set for $\tilde{B}_{i}^{P}(p,r)$. This representative set aligns better with the geometry of $B(p,r)$, and we formally define the following properties that a proper representative set should satisfy. We show in Lemma 3.7 that such proper representative set always exists. Then in the remainder of the proof, we work with this representative set instead of directly with $\tilde{B}_{i}^{P}$. While such representative set reveals the major structure of $\tilde{B}_{i}^{P}$, we note that it is only used in the analysis and the algorithm does not rely on it.
@@
Now we turn to showing that such representative sets exist. A natural attempt to define $\operatorname{rep}(\cdot,\cdot)$ such that it satisfies all three of distinct, monotone and ball-preserving properties, is to let $\operatorname{rep}_{i}(r,p):=B(p,r)\cap\operatorname{buks}_{i}(P)$, which is precisely the first half of the property Item 3. The issue is that it does not satisfy the distinctness Item 1. Our definition presented in the Lemma 3.7 below fixes this by only selecting a subset of $B(p,r)\cap\operatorname{buks}_{i}(P)$, roughly the nearest neighbors from $p$ in each bucket of $B(p,r)\cap\operatorname{buks}_{i}(P)$. This “nearest” property ensures the distinctness (Item 1). Using the geometric properties of the nearest neighbor, we can also show that such sets satisfy also properties Item 2 and Item 3.
@@
There exists $\{\operatorname{rep}_{i}(p,r):i\in[m],p\in P,r\geq 0\}$ that satisfies Definition 3.6.